The function f is linear and satisfies f(d+1)-f(d)=3 for all real numbers d. What is f(3)-f(5)? Please show your work.
Thanks in advance for your help!

Answer by richard1234(7193)

The equation "f(d+1) - f(d) = 3" already implies that the slope of f is 3, since we know the function is linear. Hence we can write f(x) = 3x + b for some real number b. Then f(3) - f(5) = (9+b) - (15+b) = -6.